OpenStudy (anonymous):

What is the probability of getting a license plate that has a repeated letter or digit if you live in a state in which license plates have one numeral followed by three letters followed by three numerals? (Round your answer to one decimal place.)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i have no clue lol

OpenStudy (anonymous):

_ _ _ _ _ _ _ so 7 slots right - how many letters in the alphabet ?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

26

OpenStudy (anonymous):

10+26+26+26+10+10+10 does this make sense

OpenStudy (anonymous):

this is all the possible license plates with this combination

OpenStudy (anonymous):

right but we are trying to find the probability of getting one with a repeated letter

OpenStudy (anonymous):

1 - P(all different) = 1 - 10*26*25*24*9*8*7/(26^3*10^4) = 0.5527 -> 0.6 --------------------

OpenStudy (anonymous):

does that make sense

OpenStudy (anonymous):

i believe so